# On Finding All Connected Maximum-Sized Common Subgraphs in Multiple Labeled Graphs

### v1

@@ Abstract
We present an exact algorithm for computing the connected Maximum Common Subgraph (MCS) across multiple graphs, where edges or vertices may additionally be labeled to account for possible atom types or bond types, a classical labeling used in molecular graphs. Our approach leverages modular product graphs and a modified Bron–Kerbosch algorithm to enumerate maximal cliques, ensuring all intermediate solutions are retained. A pruning heuristic efficiently reduces the modular product size, improving computational feasibility. Additionally, we introduce a graph ordering strategy based on graph-kernel similarity measures to optimize the search process. Our method is particularly relevant for bioinformatics and cheminformatics, where identifying conserved structural motifs in molecular graphs is crucial. Empirical results on molecular datasets demonstrate that our approach is exact, scalable and fast.
@@ 1 Introduction
Although most foundational work focused on the pairwise MCS problem—which is already NP-complete [6]—the need to compare multiple labelled graphs naturally arises in contexts like drug discovery and protein analysis. The MCS problem for more than two graph is significantly more challenging, prompting for heuristics that risk overlooking the truly maximal common subgraph. To address this gap, we propose an approach that can handle reasonably large sets of graphs and provably enumerate all optimal solutions. ***For completeness, we include a proof for the unlabeled and not necessarily connected MCS case in the Appendix. A more technical proof for the labeled and connected case will appear in an extended version of this paper.
Our proposed method is related in spirit to the MultiMCS algorithm of Hariharan et al. [7], which also applies modular products for multiple MCS A more detailed discussion of related work is given in Appendix D. Hariharan et al. proposed the following open challenge: How do we speed up the computation without losing provable correctness for the vast majority of molecules? We solve this problem by providing an exact and fast algorithm to list all (connected) maximum-sized common subgraphs in multiple labeled graphs. As the code and test data of Hariharan et al. are not publicly accessible and no formal proof of correctness is provided, it is not straightforward to compare results. Moreover, the lack of rigorous definitions permits varied interpretations, underlining the complexity of establishing correctness in multiple-MCS settings, an issue our approach seeks to address.
@@
The remainder of this paper is organized as follows. In Section 2, we provide formal definitions required throughout the paper. In Section 3, we describe how pairwise MCS can be computed using a modified Bron–Kerbosch-based approach, including the key techniques to improve runtime. We evaluate out methods in different data sets in Section 4.
@@ 2 Finding maximum common subgraphs
We consider simple undirected graphs $G=(V,E)$ with a finite vertex set $V(G)\coloneqq V$ and an edge set $E(G)\coloneqq E$. We write $G\simeq H$ if the graphs $G$ and $H$ are isomorphic. If $H$ is a subgraph of both $G_{1}$ and $G_{2}$, then it is called a common subgraph. A common subgraph is a maximum-vertex common subgraph (MVCS) of $G_{1}$ and $G_{2}$ if $|V(H)|$ is maximized among all common subgraphs of $G_{1}$ and $G_{2}$. Similarly, a maximum-edge common subgraph (MECS) $H$ is a common subgraph that maximizes $|E(H)|$. The latter definition naturally generalizes to common subgraphs of more then two graphs.
@@
The modular product can naturally be generalized to more than two factors. However, the product is not associative in general, i.e., $({G_{1}}\star{G_{2}})\star{G_{3}}\not\simeq{G_{1}}\star({G_{2}}\star{G_{3}})$ may happen. To simplify notation, we use the convention that $\star_{i=1}^{n}G_{i}\coloneqq(((G_{1}\star G_{2})\star G_{3})\dots\star G_{n})$ and that $v=(v_{1},v_{2},v_{3},\dots,v_{i},\dots,v_{n})\coloneqq(((v_{1},v_{2}),v_{3})\dots,v_{i}),\dots,v_{n})\in V(G)$. Given $G=\star_{i=1}^{n}G_{i}$, the projection onto the $i$-th factor is the map $p_{i}\colon V(G)\to V(G_{i})$, defined by $p_{i}((v_{1},v_{2},v_{3},\dots,v_{i},\dots,v_{n}))=v_{i}$. For a given subgraph $H\subseteq G$, we often write $p_{i}(H)$ for the subgraph of $G_{i}$ induced by the vertices $p_{i}(v)$ with $v\in V(H)$.
@@
For Task (1), assume that we want to compute an MVCS of the graphs $G_{1},G_{2},\dots,G_{n}$. We first compute all inclusion-maximal cliques in $G_{1}\star G_{2}$ and store them in the set $\mathcal{K}$. Then, for each $K\in\mathcal{K}$, we consider the subgraphs $H^{i}_{K}=p_{i}(K)$ in $G_{i}$, $1\leq i\leq 2$. By definition of the modular product, we have $H^{1}_{K}\simeq H^{2}_{K}$ for each $K\in\mathcal{K}$. We then collect a representative for each $K\in\mathcal{K}$, defined as $H_{K}\coloneqq H^{2}_{K}$, and store all representatives in the set $\mathcal{H}_{1,2}$. For each $H\in\mathcal{H}_{1,2}$, we repeat this process: we determine the set of all inclusion-maximal cliques in $H\star G_{3}$ and obtain the new set $\mathcal{K}$ as the collection of all inclusion-maximal cliques in $H\star G_{3}$ for all $H\in\mathcal{H}_{1,2}$. Taking the representatives $p_{3}(K)$ of all such cliques $K\in\mathcal{K}$ yields the set $\mathcal{H}_{1,2,3}$. We repeat this process recursively until we derive the set $\mathcal{H}_{1,\dots,n}$. The next two results (proofs are provide in the Appendix) show that this approach provides exact solutions for the MVCS problem and can be used to determine all MVCS of $G_{1},\dots,G_{n}$.
@@
Finally, we consider Task (4), where we may also track specific atom types (vertex labels) or bond types (edge labels). In this case, $H$ is a labeled MVCS or MECS if it is an induced subgraph of all $G_{1},\dots,G_{n}$ while preserving all edge and vertex labels. To achieve this, we extend the type-1 and type-0 classifications in the product by introducing additional types based on edge and vertex labels. Specifically, we define an edge or vertex as type-A if it has a label from a predefined set A of allowed types. For example, edge labels may distinguish between double and single hydrogen bonds, while vertex labels may represent different atom types. For non-labeled graphs, the set A defaults to $\{\texttt{1}\}$, defining type-1 and type-0 edges, while vertices remain unlabeled. In labeled graphs, connectedness refers to paths containing only edges or vertices of a particular type. To be more precise, a graph is type-A-connected if a path exists between any two vertices consisting solely of type-A edges. Note, type-A-connected clique are also knows c-clique, see [10]. If $G_{1}$ and $G_{2}$ are A-labeled graphs, their labels transfer to the product $G_{1}\star G_{2}$ by ensuring that an edge $e=\{(u,v),(u^{\prime},v^{\prime})\}$ or a vertex $x=(u,v)$ in $G_{1}\star G_{2}$ inherits label $\ell\in\texttt{A}$ if both corresponding edges $\{u,u^{\prime}\}$ and $\{v,v^{\prime}\}$, or vertices $u$ and $v$, in the factors share label $\ell$. Otherwise, $e$ or $x$ are marked, for instance, as inconsistent. Clique finding in A-labeled products follows the same approach as in unlabeled products, with a simple relabeling step: “consistent” type-A edges in the product that correspond to edges in the factors are reassigned as type-1, “consistent” type-A edges in the product that correspond to non-edges in the factors are reassigned as type-0 while all “inconsistent” edges are simply removed and do not appear as edges in the product.
@@ 3 Methods
The methods established here implicitly create a search tree, where all leaves are potential maximum common subgraphs. It turns out that a lot of branches of this search tree can be pruned by performing a depth first search and finding a candidate maximum common subgraph and then backtracking and pruning the search tree based on the current best candidate. The basic idea is as follows. To recall, $\mathcal{H}_{1,\dots,j}$ denotes the candidate sets of the MVCS of the first $j$ graph $G_{1},\dots,G_{j}$. Here, we compute first $\mathcal{H}_{1,2}$ as outlined in Section 2. Afterwards, we take one $H\in\mathcal{H}_{1,2}$ compute a candidate set $\mathcal{H}^{\prime}_{1,2}$ only for $H\star G_{3}$ which results in a subset $\mathcal{H}^{\prime}_{1,2,3}\subseteq\mathcal{H}_{1,2,3}$. In this way, we obtain a subset $\mathcal{H}^{\prime}_{1,\dots,n}\subseteq\mathcal{H}_{1,\dots,n}$. Now, we can remove all elements $H\in\mathcal{H}_{1,2}$ for which $|V(H)|<m$ where $m$ denotes the size of elements $H^{\prime}\in\mathcal{H}^{\prime}_{1,\dots,n}$ for which $|V(H^{\prime})|$ is maximum. This, in turn, can heavily reduce the size of elements in $\mathcal{H}_{1,2}$ and, thus, the number of candidates on which we need to recurse.
@@
To further optimize runtime, we refine the type-A edges in $G=H\star G_{j}$ before determining the type-A connected components $G_{1}^{\prime\prime},\dots,G_{k}^{\prime\prime}$. Specifically, we relabel edges $\{(u,v),(u^{\prime},v^{\prime})\}$ as ”inconsistent” whenever there exists no simple type-A $uv$-path in $H$ that is isomorphic to a simple type-A $u^{\prime}v^{\prime}$-path in $G_{j}$. These inconsistent edges indicate that a subgraph of $H$ containing $u$ and $u^{\prime}$ cannot be isomorphic to any subgraph of $G_{j}$ containing $v$ and $v^{\prime}$, meaning $(u,v)$ and $(u^{\prime},v^{\prime})$ cannot be part of the same type-A connected clique corresponding to a maximal type-A connected common subgraph of $H$ and $G_{j}$. When partitioning $G$ into type-A connected components $G_{1}^{\prime},\dots,G_{k}^{\prime}$, we include only type-A edges that are not marked as ”inconsistent”. We then compute maximal type-A connected cliques in $H\star G_{i}^{\prime\prime}$, ensuring they do not contain inconsistent edges. Although these edge-removal operations are based on computationally hard problems, Section 4 shows that they significantly speed up clique detection—likely due to the moderate size of the molecular graphs under consideration.
@@
As we will see in Section 4, the ordering of input graphs significantly impacts the algorithm’s runtime. Consider three graphs ordered as $[G_{1},G_{2},G_{3}]$ versus $[G_{1},G_{3},G_{2}]$. If the number of maximal common subgraphs between $G_{1}$ and $G_{2}$ is much larger than between $G_{1}$ and $G_{3}$, then $|\mathcal{H}_{12}|>|\mathcal{H}_{13}|$. Since we compute $H\star G_{3}$ for each $H\in\mathcal{H}_{12}$ and $H\star G_{2}$ for each $H\in\mathcal{H}_{13}$, the ordering $[G_{1},G_{3},G_{2}]$ results in fewer graphs requiring clique detection, leading to improved efficiency. However, determining an optimal ordering a priori is challenging, as estimating the number of maximal connected common subgraphs between each pair of graphs is itself NP-hard. To address this, we use an efficient heuristic that greedily selects graph pairs with low similarity, based on the intuition that graphs with lower similarity tend to have fewer maximal common subgraphs. Our procedure for ordering input graphs works as follows. Given a set of input graphs, $\{G_{1},\dots,G_{n}\}$, and a similarity measure $\kappa\colon\mathcal{G}\times\mathcal{G}\rightarrow\mathbb{R}$, we first select the two graphs, $G_{i}$ and $G_{j}$, which minimize $\kappa_{ij}\coloneqq\kappa(G_{i},G_{j})$ over all pairs of graphs. We let the tentative ordering be $O=[G_{i},G_{j}]$. Then we proceed as follows: given a tentative ordering $O=[G_{l_{1}},\dots,G_{l_{k}}]$ we select the graph $G_{p}\in\{G_{1},\dots,G_{n}\}-O$, which minimizes $\max\{\kappa_{l_{1}p},\dots,\kappa_{l_{k}p}\}.$ We update the tentative ordering to include $G_{p}$, i.e. $O=[G_{l_{1}},\dots,G_{l_{k}},G_{p}]$. We obtain our final ordering $O$ when $\{G_{1},\dots,G_{n}\}-O=\emptyset$.
@@
For similarity estimation, we employ two pairwise graph similarity measures. On of them is based on graph kernels [15], functions $\kappa\colon\mathcal{G}\times\mathcal{G}\rightarrow\mathbb{R}$ widely used in machine learning to approximate graph similarity. While standard graph kernels must be positive semi-definite, we do not impose this requirement. In particular, we used three of the standard graph kernels: the vertex histogram kernel(VH) [16], the Weisfeiler-Lehman optimal assignment kernel (WL) [17] and the neighborhood subgraph pairwise distance kernel (NSPD) [16]). The other similarity measure, called the “minmax” similarity, is defined as follows. We compute the minmax similarity of two graphs, $G_{i}$ and $G_{j}$, by first partitioning $G_{i}\star G_{j}$ into its type-A connected components, $G_{1}^{\prime},\dots,G_{k}^{\prime}$ as outlined in paragraph Techniques for finding type-A connected cliques. The minmax similarity of $G_{i}$ and $G_{j}$ is then computed as $\kappa_{ij}\coloneqq\max\big{\{}|V(G_{\ell}^{\prime})|\mid\ell=1,\dots,k\big{\}}$. Using the minmax similarity measure with the greedy ordering as outlined above, we greedily minimize the size of the maximum sized graph in which we have to find cliques in each of the steps. The computational results in Section 4 demonstrate a strong relationship between the chosen similarity measures and the number of maximal connected common subgraphs, validating the choice of these measures.

### v2

@@ Abstract
We present an exact algorithm for computing all common subgraphs with the maximum number of vertices across multiple graphs. Our approach is further extended to handle the connected Maximum Common Subgraph (MCS), identifying the largest common subgraph in terms of either vertices or edges across multiple graphs, where edges or vertices may additionally be labeled to account for possible atom types or bond types, a classical labeling used in molecular graphs. Our approach leverages modular product graphs and a modified Bron–Kerbosch algorithm to enumerate maximal cliques, ensuring all intermediate solutions are retained. A pruning heuristic efficiently reduces the modular product size, improving computational feasibility. Additionally, we introduce a graph ordering strategy based on graph-kernel similarity measures to optimize the search process. Our method is particularly relevant for bioinformatics and cheminformatics, where identifying conserved structural motifs in molecular graphs is crucial. Empirical results on molecular datasets demonstrate that our approach is scalable and fast.
@@ 1 Introduction
Although most foundational work focused on the pairwise MCS problem—which is already NP-complete [6]—the need to compare multiple labelled graphs naturally arises in contexts like drug discovery and protein analysis. The MCS problem for more than two graph is significantly more challenging, prompting for heuristics that risk overlooking the truly maximal common subgraph. To address this gap, we propose an approach that can handle reasonably large sets of graphs and provably enumerate all optimal solutions***For completeness, we include a proof for the unlabeled and not necessarily connected MCS case in the Appendix which is not included in the conference version but available on arXiv [7]. A more technical proof for the labeled and connected case will appear in an extended version of this paper..
Our proposed method is related in spirit to the MultiMCS algorithm of Hariharan et al. [8], which also applies modular products for multiple MCS A more detailed discussion of related work is given in Appendix D in [7]. Hariharan et al. proposed the following open challenge: How do we speed up the computation without losing provable correctness for the vast majority of molecules? We solve this problem by providing an exact and fast algorithm to list all (connected) maximum-sized common subgraphs in multiple labeled graphs. As the code and test data of Hariharan et al. are not publicly accessible and no formal proof of correctness is provided, it is not straightforward to compare results. Moreover, the lack of rigorous definitions permits varied interpretations, underlining the complexity of establishing correctness in multiple-MCS settings, an issue our approach seeks to address.
@@
The remainder of this paper is organized as follows. In Section 2, we provide formal definitions required throughout the paper. In Section 3, we describe how pairwise MCS can be computed using a modified Bron–Kerbosch-based approach, including the key techniques to improve runtime. We evaluate our methods on different data sets in Section 4.
@@ 2 Finding maximum common subgraphs
We consider simple undirected graphs $G=(V,E)$ with a finite vertex set $V(G)\coloneqq V$ and an edge set $E(G)\coloneqq E$. We write $G\simeq H$ if the graphs $G$ and $H$ are isomorphic. A “pure” subgraph of a graph $G$ is a graph $H$ satisfying $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. Moreover, $H$ is a subgraph of $G$ if $G$ contains a pure subgraph $H^{\prime}$ such that $H\simeq H^{\prime}$. If $H$ is a subgraph of both $G_{1}$ and $G_{2}$, then it is called a common subgraph. A common subgraph is a maximum-vertex common subgraph (MVCS) of $G_{1}$ and $G_{2}$ if $|V(H)|$ is maximized among all common subgraphs of $G_{1}$ and $G_{2}$. Similarly, a maximum-edge common subgraph (MECS) $H$ is a common subgraph that maximizes $|E(H)|$. The latter definition naturally generalizes to common subgraphs of more than two graphs.
@@
The modular product can naturally be generalized to more than two factors. However, the product is not associative in general, i.e. $({G_{1}}\star{G_{2}})\star{G_{3}}\not\simeq{G_{1}}\star({G_{2}}\star{G_{3}})$ may happen. To simplify notation, we use the convention that $\star_{i=1}^{n}G_{i}\coloneqq(((G_{1}\star G_{2})\star G_{3})\dots\star G_{n})$ and that $v=(v_{1},v_{2},v_{3},\dots,v_{i},\dots,v_{n})\coloneqq(((v_{1},v_{2}),v_{3})\dots,v_{i}),\dots,v_{n})\in V(\star_{i=1}^{n}G_{i})$. Given $G=\star_{i=1}^{n}G_{i}$, the projection onto the $i$-th factor is the map $p_{i}\colon V(G)\to V(G_{i})$, defined by $p_{i}((v_{1},v_{2},v_{3},\dots,v_{i},\dots,v_{n}))=v_{i}$. For a given subgraph $H\subseteq G$, we often write $p_{i}(H)$ for the subgraph of $G_{i}$ induced by the vertices $p_{i}(v)$ with $v\in V(H)$.
@@
For Task (1), assume that we want to compute an MVCS of the graphs $G_{1},G_{2},\dots,G_{n}$. We first compute all inclusion-maximal cliques in $G_{1}\star G_{2}$ and store them in the set $\mathcal{K}$. Then, for each $K\in\mathcal{K}$, we consider the subgraphs $H^{i}_{K}=p_{i}(K)$ in $G_{i}$, $1\leq i\leq 2$. By definition of the modular product, we have $H^{1}_{K}\simeq H^{2}_{K}$ for each $K\in\mathcal{K}$. We then collect a representative for each $K\in\mathcal{K}$, defined as $H_{K}\coloneqq H^{2}_{K}$, and store all representatives in the set $\mathcal{H}_{1,2}$. For each $H\in\mathcal{H}_{1,2}$, we repeat this process: we determine the set of all inclusion-maximal cliques in $H\star G_{3}$ and obtain the new set $\mathcal{K}$ as the collection of all inclusion-maximal cliques in $H\star G_{3}$ for all $H\in\mathcal{H}_{1,2}$. Taking the representatives $p_{3}(K)$ of all such cliques $K\in\mathcal{K}$ yields the set $\mathcal{H}_{1,2,3}$. We repeat this process recursively until we derive the set $\mathcal{H}_{1,\dots,n}$. The next two results (proofs are provide in [7]) show that this approach provides exact solutions for the MVCS problem and can be used to determine all MVCS of $G_{1},\dots,G_{n}$.
@@
Finally, we consider Task (4), where we may also track specific atom types (vertex labels) or bond types (edge labels). In this case, $H$ is a labeled MVCS or MECS if it is an induced subgraph of all $G_{1},\dots,G_{n}$ while preserving all edge and vertex labels. To achieve this, we extend the type-1 classifications in the product by introducing additional types based on edge and vertex labels while keeping type-0 edges unchanged. Specifically, we define an edge or vertex as type-A if it is type-1 and has, in addition, a label from a predefined set A of allowed types. For example, edge labels may distinguish between double and single hydrogen bonds, while vertex labels may represent different atom types. For non-labeled graphs, the set A defaults to $\{\texttt{1}\}$, defining type-1 and type-0 edges, while vertices remain unlabeled. In labeled graphs, connectedness refers to paths containing only edges or vertices of a particular type. To be more precise, a graph is type-A-connected if a path exists between any two vertices consisting solely of type-A edges. Note, type-A-connected clique are also knows c-clique, see [11]. If $G_{1}$ and $G_{2}$ are A-labeled graphs, their labels transfer to the product $G_{1}\star G_{2}$ by ensuring that an edge $e=\{(u,v),(u^{\prime},v^{\prime})\}$ or a vertex $x=(u,v)$ in $G_{1}\star G_{2}$ inherits label $\ell\in\texttt{A}$ if both corresponding edges $\{u,u^{\prime}\}$ and $\{v,v^{\prime}\}$, or vertices $u$ and $v$, in the factors share label $\ell$. Otherwise, $e$ or $x$ are marked, for instance, as inconsistent. Clique finding in A-labeled products follows the same approach as in unlabeled products, with a simple relabeling step: “consistent” type-A edges in the product that correspond to edges in the factors are reassigned as type-1, edges in the product that correspond to non-edges in the factors are reassigned as type-0 while all “inconsistent” edges are simply removed and do not appear as edges in the product.
@@ 3 Methods
The methods established here implicitly create a search tree, where all leaves are potential maximum common subgraphs. It turns out that a lot of branches of this search tree can be pruned by performing a depth-first search and finding a candidate maximum common subgraph and then backtracking and pruning the search tree based on the current best candidate. The basic idea is as follows. To recall, $\mathcal{H}_{1,\dots,j}$ denotes the candidate sets of the MVCS of the first $j$ graph $G_{1},\dots,G_{j}$. Here, we compute first $\mathcal{H}_{1,2}$ as outlined in Section 2. Afterwards, we take one $H\in\mathcal{H}_{1,2}$ compute a candidate set $\mathcal{H}^{\prime}_{1,2}$ only for $H\star G_{3}$ which results in a subset $\mathcal{H}^{\prime}_{1,2,3}\subseteq\mathcal{H}_{1,2,3}$. In this way, we obtain a subset $\mathcal{H}^{\prime}_{1,\dots,n}\subseteq\mathcal{H}_{1,\dots,n}$. Now, we can remove all elements $H\in\mathcal{H}_{1,2}$ for which $|V(H)|<m$ where $m$ denotes the size of elements $H^{\prime}\in\mathcal{H}^{\prime}_{1,\dots,n}$ for which $|V(H^{\prime})|$ is maximum. This, in turn, can heavily reduce the size of elements in $\mathcal{H}_{1,2}$ and, thus, the number of candidates on which we need to recurse.
@@
As outlined at the end of Section 2, for finding cliques in A-labeled products, we relabeled “consistent” type-A edges as type-1 edges, all edges in the product referring to non-edges in the factors as type-0 while all other edges in the product have been removed. To further optimize runtime, we remove additional type-0 edges in $G=H\star G_{j}$ before determining the type-A connected components $G_{1}^{\prime\prime},\dots,G_{k}^{\prime\prime}$. Specifically, we remove type-0 edges $\{(u,v),(u^{\prime},v^{\prime})\}$ from $G=H\star G_{j}$ whenever there exists no simple type-A $uv$-path in $H$ that is isomorphic to a simple type-A $u^{\prime}v^{\prime}$-path in $G_{j}$. Those particular type-0 edges $\{(u,v),(u^{\prime},v^{\prime})\}$ indicate that any type-A connected subgraph of $H$ containing $u$ and $u^{\prime}$ cannot be isomorphic to any type-A connected subgraph of $G_{j}$ containing $v$ and $v^{\prime}$, meaning $(u,v)$ and $(u^{\prime},v^{\prime})$ cannot be part of the same type-A connected clique corresponding to a maximal type-A connected common subgraph of $H$ and $G_{j}$. When partitioning $G$ into type-A connected components $G_{1}^{\prime},\dots,G_{k}^{\prime}$ (as outlined in the preceding part), we include only “consistent” type-A edges. Removal of the additional type-0 edges as explained here yields a finer partition $G_{1}^{\prime\prime},\dots,G^{\prime\prime}_{k^{\prime}}$. We then compute maximal type-A connected cliques in $H\star G_{i}^{\prime\prime}$. Although these type-0 edge-removal operations are based on computationally hard problems, Section 4 shows that they significantly speed up clique detection—likely due to the moderate size of the molecular graphs under consideration.
@@
As we will see in Section 4, the ordering of input graphs significantly impacts the algorithm’s runtime. Consider three graphs ordered as $[G_{1},G_{2},G_{3}]$ versus $[G_{1},G_{3},G_{2}]$. If the number of maximal common subgraphs between $G_{1}$ and $G_{2}$ is much larger than between $G_{1}$ and $G_{3}$, then $|\mathcal{H}_{12}|>|\mathcal{H}_{13}|$. Since we compute $H\star G_{3}$ for each $H\in\mathcal{H}_{12}$ and $H\star G_{2}$ for each $H\in\mathcal{H}_{13}$, the ordering $[G_{1},G_{3},G_{2}]$ results in fewer graphs requiring clique detection, leading to improved efficiency. However, determining an optimal ordering a priori is challenging, as determining maximal connected common subgraphs between each pair of graphs is itself NP-hard [6]. To address this, we use an efficient heuristic that greedily selects graph pairs with low similarity, based on the intuition that graphs with lower similarity tend to have fewer maximal common subgraphs. Our procedure for ordering input graphs works as follows. Given a set of input graphs, $\{G_{1},\dots,G_{n}\}$, and a similarity measure $\kappa\colon\mathcal{G}\times\mathcal{G}\rightarrow\mathbb{R}$, we first select the two graphs, $G_{i}$ and $G_{j}$, which minimize $\kappa_{ij}\coloneqq\kappa(G_{i},G_{j})$ over all pairs of graphs. We let the tentative ordering be $O=[G_{i},G_{j}]$. Then we proceed as follows: given a tentative ordering $O=[G_{l_{1}},\dots,G_{l_{k}}]$ we select the graph $G_{p}\in\{G_{1},\dots,G_{n}\}-O$, which minimizes $\max\{\kappa_{l_{1}p},\dots,\kappa_{l_{k}p}\}.$ We update the tentative ordering to include $G_{p}$, i.e. $O=[G_{l_{1}},\dots,G_{l_{k}},G_{p}]$. We obtain our final ordering $O$ when $\{G_{1},\dots,G_{n}\}-O=\emptyset$.
@@
For similarity estimation, we employ two pairwise graph similarity measures. One of them is based on graph kernels [16], functions $\kappa\colon\mathcal{G}\times\mathcal{G}\rightarrow\mathbb{R}$ widely used in machine learning. While standard graph kernels must be positive semi-definite, we do not impose this requirement. In particular, we used three of the standard graph kernels: the vertex histogram kernel(VH) [17], the Weisfeiler-Lehman optimal assignment kernel (WL) [16] and the neighborhood subgraph pairwise distance kernel (NSPD) [17]). The other similarity measure, called the “minmax” similarity, is defined as follows. We compute the minmax similarity of two graphs, $G_{i}$ and $G_{j}$, by first partitioning $G_{i}\star G_{j}$ into its type-A connected components, $G_{1}^{\prime},\dots,G_{k}^{\prime}$ as outlined in paragraph Techniques for finding type-A connected cliques. The minmax similarity of $G_{i}$ and $G_{j}$ is then computed as $\kappa_{ij}\coloneqq\max\big{\{}|V(G_{\ell}^{\prime})|\mid\ell=1,\dots,k\big{\}}$. Using the minmax similarity measure with the greedy ordering as outlined above, we greedily minimize the size of the maximum sized graph in which we have to find cliques in each of the steps. The computational results in Section 4 demonstrate a strong relationship between the chosen similarity measures and the number of maximal connected common subgraphs, validating the choice of these measures.